OpenStudy (anonymous):

The perimeter of a rectangle is 38m. Find the dimensions of the rectangle that will contain the greatest area.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

anyone?

OpenStudy (mathstudent55):

The area of a rectangle is A = LW Let the length = x, and the width = y The perimeter of a rectangle is P = 2x + 2y In this case, P = 38, so 38 = 2x + 2y, which can be solved for y: 2y = 38 - 2x y = 19 - x The area is A = LW. L = x, and W = 19 - x, so the area A = x(19 - x) A = 19x - x^2 Now differentiate with respect to x and set the derivative equal to zero to look for a maximum value: A' = 19 - 2x 19 - 2x = 0 2x = 19 x = 9.5 y = 19 - 9.5 y = 9.5 This means that L = 9.5 m and W = 9.5 m and the rectangle of greatest area is really a square with side 9.5 m.

OpenStudy (anonymous):

hmm. I see. thank you so much @mathstudent55 :)

OpenStudy (mathstudent55):

wlcm